### v2

# Fine-grained Hallucination Detection and Editing for Language Models

Abstract
Large language models (LMs) are prone to generate diverse factually incorrect statements, which are widely called hallucinations. Current approaches predominantly focus on coarse-grained automatic hallucination detection or editing, overlooking nuanced error levels. In this paper, we propose a novel task—automatic fine-grained hallucination detection—and present a comprehensive taxonomy encompassing six hierarchically defined types of hallucination. To facilitate evaluation, we introduce a new benchmark that includes fine-grained human judgments on two LM outputs across various domains. Our analysis reveals that ChatGPT and Llama2-Chat exhibit hallucinations in 60% and 75% of their outputs, respectively, and a majority of these hallucinations fall into categories that have been underexplored. As an initial step to address this, we train Fava, a retrieval-augmented LM by carefully designing synthetic data generations to detect and correct fine-grained hallucinations. On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT on fine-grained hallucination detection by a large margin though a large room for future improvement still exists. Fava’s suggested edits also improve the factuality of LM-generated text, resulting in 5-10% FActScore improvements. [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
1 Introduction
Large language models (LMs; Brown et al. 2020) can generate highly fluent and plausible text. However, these models are prone to produce hallucinations—factually incorrect generations—which often impedes their deployment in real-world applications Mallen et al. (2022). Prior work on hallucinations in natural language generation often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al. (2023). On the contrary, escalating apprehensions have been articulated with respect to LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al. (2023); Huang et al. (2022).
Figure 1: Overview of fine-grained detection system and the proposed taxonomy for hallucinations.
Such prevalence of hallucinations in LM outputs has motivated the development of automatic hallucination detection Min et al. (2023) or editing systems Gao et al. (2022); Chen et al. (2023). However, those existing systems tend to treat diverse types of factual errors uniformly, often simplifying them into binary categories such as “factual” or “not factual” with a focus on certain domains or specific types of errors. While entity-level contradictions (Figure 1 top) are usually evident and can be easily rectified with a single reference, LMs often fabricate entities that lack real-world existence (Figure 1 bottom). This variability underscores the need for a more fine-grained approach to hallucination detection for both model development and human verification, and we argue that a detection system should provide precise identification of hallucination sequences, discern different error types, and suggest potential refinements.
In this work, we introduce a new task, benchmark, and model to facilitate automatic fine-grained hallucination detection (Figure 1). We first introduce a new taxonomy of hallucinations and tasks (Section 3), and collect a high-quality benchmark to test models’ ability to provide fine-grained hallucination detection (Section 4). We then introduce a new model that is capable of providing span-level hallucination detection and editing (Section 5) as the first step towards building reliable fine-grained hallucination detection systems.
Our taxonomy (Figure 2; Table 1) hierarchically classifies factual errors in LM generations into six categories. We specifically focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters. Based on this taxonomy, we propose a new task of fine-grained hallucination detection, in which a system needs to identify locations of factual errors in a response along with fine-grained hallucination types.
We carefully annotate two widely-used LMs’ responses to diverse instruction-following and knowledge-intensive queries and construct the first fine-grained hallucination annotation benchmark. Our benchmark consists of about 400 responses of ChatGPT and Llama2-Chat 70B with span-level annotations of hallucination types based on our proposed taxonomy. Our benchmark reveals on average 1.9 and 3.4 hallucinations per response in ChatGPT and Llama2-Chat, respectively. It also shows that in addition to the widely-studied entity-level errors, other error types like unverifiable sentences or invented concepts make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
Due to the expensive costs of annotations, creating training data for this new task is challenging and our preliminary experiments find proprietary LMs such as ChatGPT perform poorly in fine-grained detection. To overcome those challenges, we design an LM-based synthetic data generation process, resulting in 35k training instances. On this training data, we fine-tune a new retrieval-augmented LM, Fava, which can identify and mark hallucinations at span level using a unified syntax. We compare Fava with state-of-the-art LMs including ChatGPT on fine-grained hallucination detection and editing in LM outputs. Fava significantly outperforms ChatGPT with and without external knowledge by 22.1% on fine-grained hallucination detection and 14.2% on binary detection. Furthermore, Fava effectively fixes factual errors in diverse LM outputs—Fava editing improves FActScore Min et al. (2023) of Alapaca 7, 13B, and ChatGPT output by 4.4%, 9.3% and 3.3%, respectively.
2 Related Work
Hallucinations in natural language generation.
Taxonomy of hallucinations in certain downstream tasks such as summarization Pagnoni et al. (2021), text simplifications Devaraj et al. (2022), or knowledge-grounded dialogues Dziri et al. (2022) have been studied. Such prior taxonomies often assume the existence of specific source text and focus on faithfulness to the source Ji et al. (2023), which is qualitatively different from LMs’ factual hallucinations grounded in world knowledge. In addition, prior taxonomies include many task-specific categories that cannot be directly adapted to describe hallucinations in LM generations.
Detecting and editing hallucinations in LMs.
Several recent works study methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al. (2023); Min et al. (2023), or focusing on entity-level factual errors editing Gao et al. (2022); Chen et al. (2023). Recent surveys have attempted to categorize hallucinations in LM generations, but such categorizations often stretch the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023). In this work, we propose a fine-grained taxonomy for fact-conflicting hallucinations in long-form text generation to improve factual consistency. While prior work often develops factuality verification systems on top of proprietary LMs Chern et al. (2023), we introduce a carefully designed training pipeline for a smaller yet more competitive fine-grained detection system. Unlike prior qualitative analysis on hallucinations in LM responses on certain knowledge-intensive queries (e.g., open-domain QA, biography generation; Liu et al., 2023b; Min et al., 2023), we annotate responses to diverse queries in multiple domains and reveal hallucination distributions vary across domains.
Figure 2: An overview of our fine-grained hallucination taxonomy.
Type Example ChatGPT Llama2
 Entity Lionel Andrés Messi was born on June 12 24, 1987. 46.9% 46.5%
 Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 2.8% 2.5%
 Contradictory Messi has yet to gain captaincy for the Argentina national football team. 12.0% 14.1%
 Invented Messi is also famous for his discovery of the famous airplane kick technique. 4.6% 23.0%
 Subjective Lionel Messi is the best soccer player in the world. 15.0% 8.8%
 Unverifiable In his free time, Messi enjoys singing songs for his family. 18.7% 5.1%
Table 1: Examples of hallucinations with their percentage in our human-annotated data. We also show the percentage of those errors in ChatGPT and LLama2-Chat-70B outputs.
Fact verification.
While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021). Thorne et al. (2018) introduce a large scale dataset based on Wikipedia documents for the training and evaluation of fact verification systems. Other datasets in information-dense domains like fact-checking against scientific articles (Wadden et al., 2020) and news documents (Wang, 2017) have been proposed. Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text. While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims as well.
3 Fine-grained Hallucination Detection
3.1 Focus and Definitions
In this work, we focus on open-ended text generation given information-seeking queries which often require factual knowledge. We define hallucinations as factually incorrect or unverifiable statements given external world knowledge. We operationalize world knowledge as documents from the web that are most relevant to the given statement/query according to a search algorithm. [^2] Errors in common sense, numerical, or logical reasoning are out of the scope of this work.
3.2 Hallucination Taxonomy
Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types. To develop a hallucination taxonomy for open-ended LM generations, we draw inspiration from error categories in prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022) and introduce new categories to describe more complex errors surfacing in LM generations. We conducted a pilot annotation with 9 NLP experts to discuss and refine our taxonomy to ensure good coverage across diverse error types.
Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination. We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
(1a)  Entity : Contradictory entity errors are a sub-category within contradictory statement errors where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
(1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship described in a statement is incorrect. While (1a) often involves errors in nouns, (1b) involves verbs, prepositions, or adjectives, and can be fixed by correcting the incorrect relation.
(1c)  Contradictory : Contradictory statement errors refer to statements that entirely contradict relevant evidence from the web. These are cases where the full sentence is refuted by the information provided in a given reference.
(2)  Invented : Invented errors refer to statements with concepts that do not exist based on world knowledge. These are cases when the LM generates a non-existent or entirely fabricated entity that doesn’t occur in any relevant evidence. This does not include fictional characters in creative work such as books or movies.
(3)  Subjective : Subjective errors refer to an expression or proposition that lacks universal validity and is often influenced by personal beliefs, feelings, opinions, or biases and hence cannot be judged as factually correct. Specifically, this represents cases where the LM generates a statement without any factual proposition grounded in relevant evidence.
(4)  Unverifiable : Unverifiable errors refer to statements that contain factual propositions but cannot be grounded in world evidence. These are cases where the LM generates statements with facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal or private matters).
While  Entity  or  Relation  are often phrase-level and can be fixed by minimal editing erroneous phrases, other error types can be an entire sentence or part of a sentence and should be removed from a response to make it factual.
3.3 Tasks and Metrics
We introduce two new tasks of identifying and editing fine-grained factual errors in LM outputs. Given an input query $x$ and a corresponding LM output $y$, our tasks require systems to identify all of the factual errors. Each error $e$ consists of $(e^{text},e^{type})$, indicating the factually incorrect text spans and their error types among our taxonomies, respectively. We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1) and editing factual errors $e^{text}$ accordingly (Task 2).
Figure 3: Hallucination type distribution of human annotations across 3 datasets and two models.
Task 1: Fine-grained hallucination detection.
In the first fine-grained error detection task, the system is expected to identify fine-grained errors in an LM output. Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$. Given an output $y$ consisting of $L$ sentences, we assume the availability of gold error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE). For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==\texttt{TRUE}]}$ (1)
$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}==\texttt{TRUE}]}$ (2)
Precision indicates how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model. For final score, we compute the F1 scores averaged over six error types as follows:
$\frac{1}{|\mathcal{E}|}\sum_{t\in\mathcal{E}}\frac{2\times\text{Prec}^{t}\times\text{Recall}^{t}}{\text{Prec}^{t}+\text{Recall}^{t}}$ (3)
Fine-grained error detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not. This is similar to FActScore Min et al. (2023), where we predict and evaluate factuality at the sentence level without extracting atomic facts.
Task 2: Hallucination editing.
As shown in Table 1, some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable. In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output. We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore or QAFactEval Fabbri et al. (2022). as follows:$f(\hat{y})-f(y)$, where $f$ indicates an evaluation function for factuality while $f(\hat{y})$ indicates edited output.
4 Benchmark
4.1 Human Annotation on Hallucinations
To facilitate model development in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are, we carefully annotate the first fine-grained hallucination detection benchmark. Our benchmark consists of multi-way fine-grained annotations on two LM responses to queries in multiple domains.
Source prompts.
To evaluate LMs’ factuality in the responses to diverse queries, we collect 150 information-seeking queries, spanning three different categories. First, we sample 50 knowledge-intensive queries from the Open Assistant Köpf et al. (2023) dataset, by prompting GPT-4 to judge whether each query from the dataset requires world knowledge or not. We, the authors further compose 50 instruction-following queries that require more complex reasoning and knowledge. Lastly, to enable more controlled evaluations and analysis, we synthetically create 50 queries that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template. See examples in each category in Appendix Table 6.
Annotation details.
We prompt ChatGPT Ouyang et al. (2022) and Llama2-Chat 70B Touvron et al. (2023) in a zero-shot manner and collect 300 responses to our diverse information-seeking queries. We recruited 15 students (five undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy. Due to the complexity of the task, we conducted 45-minute in-person and virtual training sessions with them. Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation. For each LM response, we assign two annotators. We present model-generated responses, input information-seeking queries, and the top 10 web search documents returned by Google Search given the original input queries. We ask human annotators to first mark a span, annotate a category of error, and suggest an edit to fix the errors for the  Entity  and  Relation  types. Our annotation interface and more details of annotations are available in Appendix A.
Analysis on annotated data.
Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains. 60.8% and 74.7% of the responses of ChatGPT and Llama2-Chat include at least one factual hallucination, respectively.  Entity  is the most widely recognized error type, making up 46.3% of the detected errors. Yet, there are diverse types of errors prevalent like invented statements or contradictory statements which make up 16.8% and 15.4% of the errors detected respectively. We found that the error distributions vary across different source prompts and LMs.  Invented  are more common in Llama2-Chat 70B. We observe fewer errors in Open Assistant subsets, possibly because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How tall is the Burj Khalifa?), which is often memorized by powerful LMs Mallen et al. (2022). On the other hand, the newly annotated instructions and WebNLG data show a higher percentage of errors.
Figure 4: Overview of data creation process in Fava.
Our annotators also agreed that our taxonomy covers most of the error types in LM responses. Furthermore, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in error detection across passages, with 75.1% agreement in error detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages. Appendix Table 8 shows examples of our annotations.
5 Model: Fava
Our benchmark analysis reveals the urgent need to study and detect such diverse types of hallucinations.
We introduce a new retrieval-augmented LM, Fava (FAct Vericaton with Augmentation). Fava is trained on high-quality synthetic training data, and at inference, it identifies and fixes fine-grained hallucinations, incorporating retrieved knowledge.
5.1 Overview
Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$. $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$. Subsequently, the editing model detects and if possible edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$. $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1.
While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2). Reliance on black-box proprietary API models also hurt reproducible. Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.2) and fine-tune a powerful expert LM consisting of 7 billion parameters (Section 5.3).
5.2 Training Data Creation
To train our $\mathcal{M}_{edit}$ model, we require a large number of erroneous LM outputs $y$ paired with edited versions $y^{*}$ with fine-grained error types and edits. Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al. (2022); Wang et al. (2023b); Asai et al. (2023), we introduce a new data creation method grounded on our fine-grained hallucination taxonomies. We prompt GPT-4 Achiam et al. (2023) and ChatGPT to diversify and noise a Wikipedia passage by inserting different types of hallucinations.
In particular, our data creation pipeline consists of three steps: seed passage generations, error insertions, and postprocessing. The first two processes create an input sequence with perturbed diverse error annotations, and postprocessing remaps those special tokens to generate the erroneous output $y$ and output with fine-grained detection results $y^{*}$. We use the original Wikipedia passage as well as passages retrieved by an off-the-shelf system as $\mathbf{C}$. Figure 4 shows the overview of the pipeline.
Seed passage generation.
We randomly sample a Wikipedia article $c$ and use it as our gold reference passage. To produce model-generated seed passages, we diversify the sampled article by paraphrasing and transferring the text to another genre via an LM. Then, we prompt ChatGPT to create a new genre of the text $t$, specifying to include information only found in the Wikipedia article $c$ to maintain factuality. We randomly sample one of the pre-specified genre types (e.g. blog article, question answer pair, tweet, etc.) for each passage. To further diversity in the dataset, we also add examples of question answering (QA) data. In particular, we prompt ChatGPT with two demonstrations from the NaturalQuestions Kwiatkowski et al. (2019) dataset to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles. See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 9.
Error insertion.
In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type). We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types. Therefore, we use ChatGPT and GPT-4 interchangeably for six different types. [^3] We use ChatGPT for the Type 1 and Type 3 while using GPT-4 for other types and insert factual errors one by one into our diversified text. In particular, given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retraining previously inserted errors and original text. The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution.
After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
Post processing.
We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$. At this stage, we also filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types). As a result, we have a training instance $(c,y,y^{*})$, consisting of the gold context (original Wikipedia paragraph) $c$, erroneous LM output $y$, and output with error tags and correct editing $y^{*}$. We also retrieve additional four relevant paragraphs from Wikipedia using Contriever Izacard et al. (2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$. As this process does not involve human annotations, we can easily scale up training instances under an API cost budget.
Statistics of data.
We generated a total of 35,074 training instances, 30,0074 of which are based on Wikipedia passages, and 5,000 are based on QA pairs described above. All of the error types are almost equally distributed (roughly 15% each). We present the error distributions in Appendix Table 10. We conduct a detailed human evaluation in Section 7.2. On average, 3.1 errors are inserted for the Wikipedia passage-based subsets, and 1.4 errors are inserted for the QA pairs.
5.3 Training and Inference
We train a smaller LM on the generated large-scale training data. In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using standard language modeling objective. We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$.
At inference time, we retrieve the top five documents from Wikipedia, [^4] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder. using Contriever-MSMARCO Izacard et al. (2022), and insert them together with an LM output that may include factual errors. The model identifies factual errors, marks phrases or sentences for deletion, and suggests edits for improving the factuality of the text.
6 Experiments
We evaluate Fava and state-of-the-art LMs with and without retrieval on the fine-grained hallucination detection (Section 6.1) and editing tasks (Section 6.2). We further conduct human evaluations on our best models (Section 6.3).
ChatGPT Generations Llama2-Chat 70B Generations
Editor  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
Rt+CGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
Fava (ours) 43.1 25.0 57.1 16.7 70.5 35.3 46.5 78.2 56.2 34.5 27.7 34.5 67.0 43.4 46.5 80.6
Table 2: Overall fine-grained detection performance. CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
6.1 Experiments for Hallucination Detection
Evaluation data.
We use our new benchmark, consisting of 364 annotated passages, [^5] In total we collected 471 annotated passages including our experimental batches, and we excluded earlier batches for evaluations. as our test data. We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.3.
Baselines.
For fine-grained hallucination detection tasks, we test three baselines. ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations. Rt-ChatGPT uses the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt. [^6] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
6.2 Experiments for Editing
Evaluation data and function.
For editing evaluations, we use two open-ended text generation data and evaluate models’ factuality based on two metrics designed to measure the factuality of text. FActScore Min et al. (2023) prompts ChatGPT and GPT 3-davinci 003 to decompose a response into a set of atomic facts and verify factuality for each of them using passages from a designated Wikipedia article. We use the original biography generation task proposed with FActScore for evaluation.
Baselines
For editing experiments, we use ChatGPT and Rt-ChatGPT, as well as Llama2-Chat 13B with and without retrieval (Llama and Rt-Llama, respectively). As in the fine-grained error detection task, we prompt the models with carefully designed prompts and two demonstrations. For all baselines, we retrieve five documents consisting of one document retrieved by string matching based on the entity name and four documents retrieved by Contriever, which are reranked at test time, using a small cross-encoder. [^7] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2
6.3 Human Evaluations
Our automatic evaluations may not fully capture the models’ abilities to detect annotated data and editing errors due to the potential subjectivity of annotations and may be affected by the performance of factuality evaluation metrics. We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT. We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents. This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at the individual detection level. Due to the cost of evaluating the factuality of all verification-worthy statements Min et al. (2023), we do not evaluate recall.
7 Results and Analysis
7.1 Results
Detection results.
Table 2 shows the fine-grained detection accuracy of Fava and baselines. We provide the full precision and recall scores in Appendix Tables 12 and 13, respectively. We also provide binary prediction performance.
Table 2 shows that while ChatGPT-based baselines show relatively good binary predictions of existence of factual errors, their fine-grained error type detection performance is significantly lower. Fava shows its strong capabilities of conducting fine-grained error detection, resulting in significant performance improvements from ChatGPT or Ret-ChatGPT on both fine-grained error detection and binary error detection. Specifically, Fava shows high accuracy on error types such as  Contradictory ,  Subjective ,  Entity . On the other hand, its performance on  Invented and  Unverifiable are still limited. Those two error types often require intensive search over many web documents beyond top few passages, while Fava by default only considers top five documents. Our inter-annotator agreements also show that there are more disagreements in those two categories, which might also deflate the performance.
ChatGPT Al-13B Al-7B
No Edit 66.7 42.5 38.8
[…]
Rt-ChatGPT 62.7 -4.0 43.9 +1.4 39.2 +0.4
Llama 52.6 -14.1 22.7 -19.8 18.6 -20.2
Rt-Llama 58.7 -8.0 48.6 +6.1 32.2 -6.6
[…]
Table 3: Overall results on FActScore. Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
Editing results.
Table 3 shows the results on FactScore. Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model size. Retrieval-augmented ChatGPT shows small improvements by editing Alpaca outputs, editing its output with retrieved documents hurts the FActScore. We found that ChatGPT is often confused with different error types and exhibits tendencies of over-predicting contradictory or subjective statements. Among other models, our methods show the largest gains in terms of factuality, showing the effectiveness of editing, in addition to its strong capability of providing fine-grained error types.
Human evaluation results.
Our human evaluation results are shown in Table 4. As in automatic evaluations, our method shows significantly better performance than Rt+ChatGPT on both editing and detection and recognizes more errors than retrieval-augmented ChatGPT. These results further demonstrate the strong capabilities of Fava detecting factual errors in LM outputs.
Model avg. $|\mathbf{E}|$ Detect (%) Edit (%)
Rt+ChatGPT 1.9 23.9 17.2
Fava (ours) 2.4 55.7 63.5
Table 4: Human evaluation results. We show the average number of detected errors and the correctness (%) of the fine-grained types and edits.
7.2 Analysis
Figure 5: Detection scores across different training instance values.
We conduct a set of ablation studies of different aspects of training data creation and inference time design choices, to understand the key factors to build a reliable fine-grained error detection system.
Effects of scale.
We evaluate our model trained on varying numbers of model-generated training data. In particular, we train three Fava variants trained on 10k, 20k, and 30k training instances, and evaluate their performance on fine-grained error detection tasks. Figure 5 shows the results of each variant and demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
Effect of retrieval.
We assess Fava’s editing performance on FActScore on Alpaca 13B outputs through various retrieval methods during inference. Specifically, we examine four distinct retrieval approaches: retrieving the top 1 document; retrieving the top 5 documents; reordering the top 5 documents; retrieving the top 4 documents and mixing them with an introductory paragraph of the target entity (entity matching). We provide the details of the ablation studies in Appendix Section C.4.
Settings FActScore
Alpaca 13B (No Edit) 42.5
Fava (Top 1) 44.2 +1.7
Fava (Top 5) 47.0 +4.5
Fava (Top 5) + reranked 47.7 +5.2
Fava (Top 5) + entity matching 50.1 +7.6
Table 5: FActScore editing results with different retrieval settings.
Table 5 shows the experimental results. Retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%. Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effects of context locations, as reported by prior work Liu et al. (2023a). Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations. Prior work reports that neural retrieval systems can struggle in rare entities Mallen et al. (2022), where term-based retrievers or simple string matching shows more robust performance. FActScore bio consists of extremely low entities, where simple entity matching may help. These results indicate although Fava demonstrates strong capabilities, there’s room for improvements in the retrieval components and how to incorporate many references.
Human evaluation of generated data.
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. We evaluate a generated instance from two aspects: (1) validity—whether the model edits do not violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples. We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
Our analysis revealed that the data generated by our system with iterative insertion obtained an average score of 1.66 for validity assessment and 1.36 for quality assessment. Meanwhile, data generated by one-shot prompting scored an average of 1.1 for validity and 0.9 for quality. Our human evaluation reveals that our LM-generated edits are indeed of higher quality than the one-shot data creation, and provide realistic factual errors.
8 Conclusions and Discussions
This work introduces a new task of automatic hallucination detection, built upon our newly introduced taxonomy, which hierarchically classifies hallucinations in LMs into six categories. We collect the first human-annotated fine-grained hallucination benchmark, consisting of span-level hallucination types, hallucination text, and edits for two LM outputs across multiple domains. As the first step, we introduce a new retrieval-augmented LM trained on synthetically generated fine-grained error editing data following our taxonomies. Empirical results show that Fava significantly outperforms strong baselines by a large margin on both editing and fine-grained detection tasks, while still large room for improvements for automated fine-grained error detection and editing.
Acknowledgement
We thank Sewon Min for fruitful discussions in the early stages. Akari Asai is supported by the IBM Fellowship. We thank Stability AI for providing computing to train and evaluate the LMs in this work, and Microsoft Accelerate Foundation Models Research Program for the access to Azure and OpenAI APIs. We thank our annotators who provide high-quality annotations for this work. This work was funded in part by the DARPA MCS program through NIWC Pacific (N66001-19-2-4031), NSF IIS-2044660, and gifts from AI2.
Appendix
Appendix A Details of Human Annotations
Seed query selections.
Table 6 shows examples of the seed queries for our initial dataset creation.
Dataset Example
WebNLG Explain A.C. Cesena, including information about ground, league.
Instruction-following Explain the differences between New York cheesecakes and Basque cheesecakes in detail.
Open-Assistant Can you tell me how tall the burj khalifa is?
Table 6: Examples of source prompts.
Annotation interface.
Figure 6 shows the interface of our annotations.
Figure 6: Annotation interface.
Details of annotations.
We hired 5 undergrads studying computer science to complete annotations. Each annotator was assigned 60 passages to annotate and had to undergo a 45 minute one on one training session to understand the task and how to navigate the annotation platform. The training session covered an in depth explanation of the six different hallucination types in our taxonomy, included a walk through of how to annotate a passage, and allowed time for annotators to ask any questions.
Examples of annotated results.
Table 7 shows the fine grained hallucination distribution across our annotations. Table 8 shows list of annotated examples from each of our annotation datasets - WebNLG, OpenAssistant, and Instruction Following dataset.
Type Percentage
Entity Errors 28.4%
Relation Errors 2%
Contradictory Statements 6.1%
Invented Statements 3%
Subjective Statements 20.2%
Unverifiable Statements 40.3%
Table 7: Hallucination percentages in our human-annotated data.
Dataset
Annotated Passage Example
WebNLG
A Long Long Way is a novel written by Sebastian Barry. It was first published in 2005 by <entity>Faber and Faber viking press</entity>, a <entity>British American</entity> publishing company. The novel tells the story of Willie Dunne, an Irish soldier who has a strong sense of loyalty to his country, but also feels torn between his Irishness and his loyalty to the British army during World War I. The novel was critically acclaimed and received several literary awards, <contradictory>including the 2005 Dublin Literary Award </contradictory> and was shortlisted for the Man Booker Prize in the same year.
OpenAssistant
<subjective>The most luxurious airline flying within the continental United States is generally regarded as JetBlue Airways</subjective>, particularly for its Mint business class service. Mint offers a private suite experience, featuring fully lie-flat seats, ample legroom and storage, premium food and beverage options, access to onboard WiFi and live TV streaming, and turn-down service complete with duvets and pillows.
Instruction Following
Some of Afghanistan’s secondary cricket home grounds include:
1.<contradictory>Kandahar Cricket Stadium, Kandahar </contradictory>
2.<invented>Mis Ainak Cricket Ground, Logar</invented>
3.<invented>Bost University Cricket Ground, Helmand</invented>
Table 8: Examples from our human annotations for each dataset.
Appendix B More Details of Synthetic Data Creation and Training
Text diversification results.
Table 9 shows the diversified text results.
Seed Passage Diversification Examples
Genre Topic
Output
Article Community Reinvestment Act
The Community Reinvestment Act (CRA) is a federal law that encourages commercial banks and savings associations to help meet the needs of borrowers in low- and moderate-income neighborhoods. The Act was passed in 1977 by Congress to reduce discriminatory credit practices against low-income neighborhoods.
Essay National Dodgeball League
The National Dodgeball League is the only professional dodgeball league in the US, and was founded in 2004. It is headed by Commissioner Edward Prentiss out of Hopkins, MN and is composed of 24 professional teams that are divided into the National and American Dodgeball Conferences. In the past, the tournament has featured teams from Canada, Denmark, UK, Japan, New Zealand and Australia.
Blog Post Red Channels
In the beginning of the 1950s, the "Counterattack" journal published a document, "Red Channels," which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association. Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history. "Red Channels" effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
Social Media Post Sandman
Check out The Sandman - a semi-retired American professional wrestler! He is best known for his career with Extreme Championship Wrestling (ECW), where he held the ECW World Heavyweight Championship a record five times. He also had stints in World Championship Wrestling (WCW), Total Nonstop Action Wrestling (TNA), and World Wrestling Entertainment (WWE). #Wrestling #ECW #WCW #TNA #WWE
News Report I Don’t Mind (James Brown song)
James Brown’s "I Don’t Mind" has made its way up the R&B "Billboard" charts, hitting number four. The track was released in 1961 and also reached number 47 on the Pop "Billboard" charts. Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
Summary Robert Lui
Robert Lui was born on February 23, 1990, and is a professional rugby league footballer from Australia. He plays either as a halfback or five-eighth for the Townsville Blackhawks in the Queensland Cup.
Speech George Sperling
Dear fellow Americans, I am here to propose a solution to improve American Sign Language communication. George Sperling suggests that with a sevenfold reduction in the bandwidth for video transmission, we can achieve this. He even argued that the telephone was originally created for the hearing impaired but it became popularized by the hearing community. Let us not forget our roots and make a change for the better.
Presentation Intro Blissful Ignorance Effect
Have you ever wondered why sometimes people who know less about a product seem to enjoy it more than those who have researched it thoroughly? This phenomenon is called the Blissful Ignorance Effect and it’s a fascinating topic in consumer behavior studies. Our presentation today will explore this effect and why it happens.
Brochure Nidulariaceae Fungi
Explore the fascinating world of Nidulariaceae fungi! This family, found in most ecological regions, includes five different genera, each with its own unique characteristics. With their tiny egg-filled structures, these fungi are a wonder to behold!
Text Message Ashes and Diamonds
Just found out about this book called Ashes and Diamonds by Jerzy Andrzejewski. It’s set during the last few days of WWII. The main character, Maciek, has to kill a Communist soldier. Sounds intense!
Table 9: Text diversification prompts. Instructions for diversification follow the following format: "Given a passage, create a(n) [genre] of 3-6 sentences using only the information present in the passage. Do not include any new information not presented in the passage. Passage: [sampled wikipedia paragraph]"
Error distribution.
Table 10 shows the distribution of error types across all our generated passages.
Percentage
 Entity 21.2%
 Relation 19.9%
 Contradictory 15.3%
 Invented 14.6%
 Subjective 14.1%
 Unverifiable 14.9%
Table 10: Statistics of generated training data
Training Details.
Our base model is Llama2-Chat 7b trained using 4xA40 GPUs. Our training code is based off Open-Instruct Wang et al. (2023a) [^8] https://github.com/allenai/open-instruct/blob/main/scripts/finetune_with_accelerate.sh. Table 11 shows the training hyperparameters.
Precision Epochs Weight Decay Warmup Ratio Learning Rate Max. Seq. Length Batch Size
BFloat16 2 0 0.03 2e-5 2048 128
Table 11: Training hyperparameters.
Appendix C More Results and Analysis
C.1 Detection Results
Overall precision and recall.
Tables 12 and 13 report the precision and recall on our curated datasets.
ChatGPT Generations Llama2-Chat 70B Generations
Model  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 12.1 25.0 50.0 6.7 7.7 0.0 13.0 35.1 17.5 13.2 36.4 15.5 18.8 9.5 19.3 59.2
R+CGPT 19.5 12.5 20.6 3.3 28.6 11.9 17.2 49.7 25.2 16.7 18.8 25.7 14.9 3.2 21.0 60.0
Ours 35.4 20.0 46.2 10.0 75.0 30.0 40.1 69.1 56.8 31.3 23.7 39.6 70.6 46.4 46.1 80.0
Table 12: Fine-grained detection task (Precision). CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
ChatGPT Generations Llama2-Chat 70B
Model  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 51.3 33.3 33.3 50.0 7.7 0.0 33.6 87.4 41.8 19.2 21.1 8.5 16.7 20.0 32.4 81.1
R+CGPT 49.9 41.7 33.3 14.3 55.6 22.4 41.7 93.1 50.9 45.8 33.3 19.6 22.2 8.8 41.0 92.5
Ours 55.2 33.3 75.0 50.0 66.7 42.9 55.5 90.0 55.5 38.5 33.3 30.6 63.7 40.6 46.8 81.2
Table 13: Fine-grained detection task (Recall). CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
C.2 Details of Human Evaluation on Model Outputs
We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT. We anonymize the model results to avoid potential biases during evaluations. Annotators count how many of the detection and editing results are correct given extensive web search.
C.3 Manual Analysis on Generated Data for Fava training
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. Prior work introduces entity-centric synthetic data creation Longpre et al. (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed. Therefore, we evaluate not only validity but also quality Table  14 shows examples of generated erroneous data from our synthetic data creation pipeline. We ask human annotators to score each category for each passage either 0, 1, or 2. These scores were defined as follows: a score of 0 indicated that less than half of the inserted errors aligned with the given criteria, a score of 1 signified that at least half of the errors met the established criteria, and a score of 2 indicated that all inserted errors met the criteria. As a comparison, we also generate training data with one-shot generation.
Original Passage
Automatic Data Generations for Training
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive Multiplayer in 1994, the release was ultimately canceled and never reached the market.
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. <invented>The game won the Shooting Star Award in 2020 for its innovative concept.</invented> Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive <entity><delete>Multiplayer</delete><mark>Singleplayer </mark></entity> in 1994, the release was ultimately canceled and never reached the market. <subjective>The game is a masterpiece that revolutionized the arcade industry and will always be remembered as a classic.</subjective>
The knapsack problem is a problem in combinatorial optimization where the goal is to determine the optimal selection of items to maximize the total value within a given weight constraint.
The knapsack problem is a problem in <entity><delete>combinatorial optimization</delete><mark>sequential search</mark></entity> where the goal is to determine the optimal selection of items to maximize the total <entity><delete>value</delete><mark>weight</mark></entity> within a given weight constraint.
Table 14: Examples of generated data from automatic data creation.
C.4 More Ablations
Retrieval ablations.
Settings FActScore
Alpaca 13B (no edits) 42.5
Fava (Top 1) 44.2 +1.7
Fava (Top 5) 47.0 +4.5
Fava (Top 5) + reranked 47.7 +5.2
+ entity search 50.1 +7.6
+ order 51.8 +9.3
Table 15: FActScore editing results with different retrieval settings.
Table 15 shows the full results of the retrieval ablations. We enhance our retrieval process by strategically reordering the retrieval results. This involves rearranging the top 5 passages from our reranked Contriever documents and entity matching document such that the documents with the highest relevance are positioned closest to the text requiring verification. Reordering passages gives 1.7% improvements, further emphasizing the importance of retrieval and careful pipeline design.